OpenStudy (anonymous):

Hey! How do you find the average value of f(x)=4x-x^2 on the interval [0,4]?

OpenStudy (anonymous):

so basically you you know your average value formula is (1/b-a) then integral of f(x) between a and b you are just using regular integration using the fundamental theorem of calculus but all your doing is multiply it by 1/b-a

OpenStudy (anonymous):

so in answer to your question is 1/4[ 2x^2 - x^3/3] integrating that between 0 and 4 is 1/4[32-64/3] = 8 -16/3 = 8/3

OpenStudy (anonymous):

ohhok..thank you!

OpenStudy (anonymous):

no problemm